# Unifying the Perspectives of NLP and Software Engineering: A Survey on Language Models for Code

### v3

@@ 3.1.3 Code-to-Text
Code-to-text tasks take code as input, and output text. Related methods are listed in Figure 7.
@@ 3.1.4 Code-to-Pattern
Code-to-pattern tasks conduct classification on code. Related methods are listed in Figure 8.
@@
- Defect detection predicts whether the input code is buggy or not, and is a standard single-sentence classification task. Nong et al. (2023); Steenhoek et al. (2023); Bi et al. (2023); Harzevili et al. (2023) provide surveys on this task.
- Clone detection predicts whether or not two pieces of code are clones of each other. In software engineering there exist four types of code clones, and the most challenging type to identify is semantic clones, i.e. syntactically dissimilar code that have the same functionality. As this task can be viewed as a two-sentence classification task, BERT-style language models have been widely applied to it. Svajlenko & Roy (2020) and Zhang & Sakurai (2021) provide comprehensive reviews on non-deep-learning based methonds for this task.
@@
- Machine code detection is another recently introduced task and aims to predict whether the input code is written by human or generated by machines. Detection of AI generated content (AIGC) has undergone thorough studies in general or specialized domains such as education (Mitchell et al., 2023; Liu et al., 2023j), but it is yet to be examined at scale on code.
@@ 3.1.5 Text-to-Text
Text-to-text tasks take text as input, and output text. Related methods are listed in Figure 8.
@@
- Log analysis aims to analyze the system logs produced by software products, for example parsing logs into structured templates or finding anomalies from raw logs. Zhu et al. (2019) provide a survey on traditional methods for this task up to 2018, and Chen et al. (2021d) give an empirical comparison between neural network based methods. Zhang et al. (2023d) also cover more recent methods for log parsing, while Landauer et al. (2022) survey methods for anomaly detection in logs.
@@ 3.1.6 NLP Point-of-View
Code completion and code infilling are also generation tasks, and correspond exactly to the two pretraining objectives given in (1) and (3), except that for code infilling only one span in the input is masked. Similarly, cloze test is an understanding task in the same form as (2).
Defect detection, clone detection, code classification, and author identification are sequence classification tasks. In these tasks, a set of labels is defined over the input, and each instance is assigned a label (e.g. for defect detection , while for author identification a possible is {Alice, Bob, John, others}). The model is then tasked to maximize
@@ 3.2 Evaluation Metrics
Of the tasks mentioned in Section 3.1, the understanding tasks are similar in form to natural language understanding tasks (Wang et al., 2018a; 2019) and evaluated likewise by metrics such as accuracy, F1 and Mean Reciprocal Rank (MRR), while short generation tasks such as identifier prediction is also evaluated by accuracy of exact matches. Code-to-text tasks are evaluated with common metrics for text generation such as BLEU (Papineni et al., 2002),
@@ 3.3 Program Synthesis
We identify four changes in program synthesis over the years: shift of coding paradigms (from example-based to intention-based), generalization in languages (from domain-specific languages to general-purpose languages), simplification of model architectures (from grammar-guided decoders to general-purpose language models), and application of execution-based feedback.
@@
Many of the early methods for program synthesis are exampled-based (Menon et al., 2013), which means they induce programs from input-output examples, often in domain-specific languages (DSLs) such as FlashFill (Devlin et al., 2017a) and Karel444FlashFill is used in Microsoft Excel for string transformation. Karel is a simple programming language for educational purpose. (Bunel et al., 2018), as these languages are usually simple in syntax and structure.
@@
As code generation models became more capable over the years, researchers started to pay attention to program synthesis in general-purpose programming languages as well. Hearthstone (Ling et al., 2016) and CONCODE (Iyer et al., 2018) are two of the early datasets, representing Python and Java respectively. Each example in Hearthstone is the description of a card in the game and its corresponding class implementation, while examples in CONCODE are simply Java methods paired with their natural-language documentation crawled from public GitHub repositories. Synthesizing programs from their corresponding natural language descriptions has since then become a standard practice in program synthesis, and has led to some of the most widely used benchmarks, such as HumanEval (Chen et al., 2021b), which has even been translated into multiple languages (Cassano et al., 2023a; Zheng et al., 2023a; Muennighoff et al., 2023). Some recent benchmarks use general-purpose languages but focus on specific domains, such as data science (Bavishi et al., 2019; Lai et al., 2023) or Jupyter notebooks (Agashe et al., 2019), while several math reasoning benchmarks have also been converted to programming tasks, including MathQA-Python (Amini et al., 2019; Austin et al., 2021) and GSM8K-Python (Cobbe et al., 2021; Chowdhery et al., 2022; Wang et al., 2023d).
@@
Many early works argue that simply treating program synthesis as a text generation task does not utilize the underlying syntax of programming languages, and thus often use syntax-enhanced decoders to inject the target syntax as prior knowledge (Yin & Neubig, 2017). LLMs, however, have demonstrated that pretrained language models are capable of generating syntactically correct programs without loss of generality. Under this setting, researches start to execute the generated programs and provide feedback to the generation model to inject the prior knowledge of code instead. This has recently led to the popularity of interactive coding, which we discuss in more detail in Section 7.1.
3.4 Repository-Level Evaluation
@@
Most evaluation tasks discussed in Section 3.1 are limited to a single file or even a single function, as cross-file code modeling poses challenges that are beyond the capability of most existing language models. Recently, however, position interpolation techniques (Chen et al., 2023b; Rozière et al., 2023; Peng et al., 2023a) have extended the context window of LLMs to hundreds of thousands of tokens, making it possible to contextualize the evaluation of code modeling within entire repositories. Several works (Shrivastava et al., 2023b; Ding et al., 2022b; Zhang et al., 2023b; Shrivastava et al., 2023a) have studied code completion leveraging repository-level context, and Liu et al. (2023i) propose RepoBench to evaluate such systems. More recently, Bairi et al. (2023) investigate the more challenging tasks of repository-level API migration and temporal editing, and Jimenez et al. (2023) introduce a corresponding benchmark, SWE-bench.
@@ 4 General Language Models for Code
Since language models scaled to hundreds of billions of parameters (Brown et al., 2020; Chowdhery et al., 2022), many of them have demonstrated non-trivial coding capability, even if they are not specifically designed or trained for code. Pioneered by Codex, researchers have also found continual pretraining on code to significantly benefit language models’ performance on code555While some works refer to this process as “finetuning on code”, it is still self-supervised in nature. Thus we choose to adopt the term “extra/additional/continual pretraining” in this work to avoid confusion with supervised in-task finetuning or instruction finetuning..
@@ 4.1 Off-the-Shelf Language Models
Closed-source models, on the other hand, perform generally better. LaMDA (Thoppilan et al., 2022) and PaLM (Chowdhery et al., 2022), whose pretraining dataset contains 12.5% and 5% code respectively, achieve 14.0 and 26.2 pass@ performance on HumanEval, while GPT-4 (OpenAI, 2023) set a staggering record of 67.0 (and an early version is reported by Bubeck et al. (2023) to be 82) that until recently has remained higher than any specialized models pretrained or instruction-finetuned for code.
| HumanEval (0) | MBPP (3) | |||
| k=1 | k=100 | k=1 | k=80 | |
| GPT-J | 11.6 | 27.7 | ||
| LaMDA | 14.0 | 47.3 | 14.8 | 62.4 |
| PaLM | 26.2 | 76.2 | 36.8 | 75.0 |
| GPT-NeoX | 15.4 | 41.2 | ||
| BLOOM | 15.5 | 55.5 | ||
| LLaMA | 23.7 | 79.3 | 37.7 | 76.8 |
| GPT-4 | 67.0/82 | |||
| LLaMA 2 | 29.9 | 89.0 | 45.0 | 81.5 |
| Phi-1.5 | 41.4 | 43.5 | ||
| Baichuan 2 | 17.1 | 30.2 | ||
| Qwen | 32.3 | 40.8 | ||
| Codex | 28.8 | 72.3 | ||
| PaLM-Coder | 36.0 | 88.4 | 47.0 | 80.8 |
| PaLM 2-S* | 37.6 | 88.4 | 50.0 | 86.8 |
| Code LLaMA | 53.7 | 94.7 | 56.2 | |
| CodeFuse | 74.4 | 61.0 | ||
More recently, the general trend has been to train smaller models with larger datasets, following the revised scaling law (Hoffmann et al., 2022). Baichuan 2 (Yang et al., 2023a), for example, is a 13B model trained on 2.6T tokens, while Qwen (Bai et al., 2023) is a 14B model trained on 3T tokens. They achieve 17.1 and 32.3 pass@1 on HumanEval, respectively. Li et al. (2023h), however, demonstrate that models as small as 1.3B can acquire coding capability that’s comparable to much larger models while also maintaining a reasonable performance on general text processing and even manifesting some emergent abilities (Wei et al., 2022c) such as chain-of-though reasoning (Wei et al., 2022d). Their model, Phi-1.5, is trained on 21B tokens of textbook data generated by ChatGPT, and 100B tokens of filtered web data from Stack Overflow and Refined Web (Penedo et al., 2023), and attains 41.4 pass@ performance on HumanEval. The exact performance of these models are presented in Table 1.

### v5

@@ 2.1.3 Code-to-Text
Code-to-text tasks take code as input, and output text. Related works are listed in Figure 6.
@@ 2.1.4 Code-to-Pattern
Code-to-pattern tasks conduct classification on code. Related works are listed in Figure 7.
@@
- Defect detection, or vulnerability detection, predicts whether the input code is buggy or not, and is a standard single-sentence classification task. Nong et al. (2023); Steenhoek et al. (2023a); Bi et al. (2023); Harzevili et al. (2023); Zhou et al. (2024) provide surveys on this task.
- Malware detection is a similar task to defect detection. However, malware differs from other types of vulnerable code in that the bugs therein are malicious - i.e. they are intentionally injected.
- Clone detection predicts whether or not two pieces of code are clones of each other. In software engineering there exist four types of code clones, and the most challenging type to identify is semantic clones, i.e. syntactically dissimilar code that have the same functionality. As this task can be viewed as a two-sentence classification task, BERT-style language models have been widely applied to it. Svajlenko & Roy (2020) and Zhang & Sakurai (2021) provide comprehensive reviews on non-deep-learning based methods for this task.
@@
- Machine code detection is another recently introduced task and aims to predict whether the input code is written by human or generated by machines. Detection of AI generated content (AIGC) has undergone thorough studies in general or specialized domains such as education (Mitchell et al., 2023; Liu et al., 2023k), but it is yet to be examined at scale on code.
@@ 2.1.5 Text-to-Text
Text-to-text tasks take text as input, and output text. Related works are listed in Figure 7.
@@
- Log analysis aims to analyze the system logs produced by software products, for example parsing logs into structured templates or finding anomalies from raw logs. Zhu et al. (2019) provide a survey on traditional methods for this task up to 2018, and Chen et al. (2021d) give an empirical comparison between neural network based methods. Zhang et al. (2023e) also cover more recent methods for log parsing, while Landauer et al. (2022) survey methods for anomaly detection in logs.
@@ 2.1.6 NLP Point-of-View
Unlike software engineering, evaluation tasks in NLP are generally categorized into understanding and generation. The former, represented by GLUE (Wang et al., 2018a) and SuperGLUE (Wang et al., 2019), emphasizes the comprehension of input text, and is typically formalized as classification, regression, sequence tagging, or span extraction. The later, on the other hand, involves autoregressive generation of text, such as machine translation and summarization.
@@
Code completion and code infilling are also generation tasks, but differ from sequence-to-sequence tasks where the input and output are related by different sequences. In these two tasks, the target is a continuation or infill of the input. They correlate closely to the language modeling objectives given in Equation (3) and (5). Similarly, cloze test takes the same form as Equation (4) but is usually considered an understanding task, as its output is usually a single token and does not involve autoregressive generation.
Defect detection, malware detection, clone detection, code classification, and author identification are sequence classification tasks. In these tasks, a set of labels is defined over the input, and each instance is assigned a label (e.g. for defect detection , while for author identification a possible is {Alice, Bob, John, others}). The model is then tasked to maximize
@@
2.1.7 Code LLMs for Low-Resource, Low-Level, and Domain-Specific Languages
In NLP, human languages are categorized into high-, middle-, and low-resource languages based on the amount of available data in each language. High-resource languages such as English are extensively studied, while low-resource languages such as Swahili and Yoruba often rely on transfer learning from other languages to improve performance due to data scarcity (Conneau et al., 2020; Xue et al., 2021; Zhang et al., 2023g).
This phenomenon also exists in code modeling, as most works listed in Figure 4-7 target the most popular programming languages such as C, Java, and Python. However, a major difference between NLP and SE is that in NLP, low-resource languages are often spoken by few people or even endangered, while in SE low-resource languages are often designed for specific domains and purposes, and thus have an active user community. Verilog, a hardware description language, is one such example, for which code modeling research is quite abundant (Thakur et al., 2023a; b; Lu et al., 2023b; Liu et al., 2023g; Tsai et al., 2023; Thorat et al., 2023; Liu et al., 2023h; Nadimi & Zheng, 2024).
Similar to NLP, many works in code modeling have studied the transfer of model capability across languages. Chen et al. (2022), for example, investigate the performance of multilingual language models on Ruby (which is one of the six languages in the popular CodeSearchNet pertaining corpus (Husain et al., 2019), but relatively low-resourced compared with the other five), finding that multilingual models have lower performance-to-time ratio compared with monolingual ones. Cassano et al. (2023a), on the other hand, propose MutiPL-T, an approach for translating high-resource training data into low-resource languages for fine-tuning. However, such research is yet scarce compared with the NLP community, but with the recent advent of colossal, multilingual datasets such as The Stack v2 (Lozhkov et al., 2024) we expect to see more of it in the future.
@@ 2.2 Evaluation Metrics
Of the tasks mentioned in Section 2.1, the understanding tasks are similar in form to natural language understanding tasks (Wang et al., 2018a; 2019) and evaluated likewise by metrics such as accuracy, F1 and Mean Reciprocal Rank (MRR), while short generation tasks such as identifier prediction is also evaluated by accuracy of exact matches. Code-to-text tasks are evaluated with common metrics for text generation such as BLEU (Papineni et al., 2002).
@@ 2.3 Program Synthesis
While dozens of evaluation tasks exist in software engineering, they have generally stayed out of the focus of the NLP community until very recently. The only exception is program synthesis, which has become a standard evaluation task for LLMs since the advent of HumanEval in 2021. Looking back at this task, we identify four changes in program synthesis over the years: shift of coding paradigms (from example-based to intention-based), generalization in languages (from domain-specific languages to general-purpose languages), simplification of model architectures (from grammar-guided decoders to general-purpose language models), and application of execution-based feedback.
@@
Many of the early methods for program synthesis are example-based (Menon et al., 2013), which means they induce programs from input-output examples, often in domain-specific languages (DSLs) such as FlashFill (Devlin et al., 2017a) and Karel222FlashFill is used in Microsoft Excel for string transformation. Karel is a simple programming language for educational purpose. (Bunel et al., 2018), as these languages are usually simple in syntax and structure.
@@
As code generation models became more capable over the years, researchers started to pay attention to program synthesis in general-purpose programming languages as well. Hearthstone (Ling et al., 2016) and CONCODE (Iyer et al., 2018) are two of the early datasets, representing Python and Java respectively. Each example in Hearthstone is the description of a card in the game and its corresponding class implementation, while examples in CONCODE are simply Java methods paired with their natural-language documentation crawled from public GitHub repositories. Synthesizing programs from their corresponding natural language descriptions has since then become a standard practice in program synthesis, and has led to some of the most widely used benchmarks, such as HumanEval (Chen et al., 2021b), which has even been translated into multiple languages (Cassano et al., 2023b; Zheng et al., 2023a; Muennighoff et al., 2024). Some recent benchmarks use general-purpose languages but focus on specific domains, such as data science (Bavishi et al., 2019; Lai et al., 2023) or Jupyter notebooks (Agashe et al., 2019), while several math reasoning benchmarks have also been converted to programming tasks, including MathQA-Python (Amini et al., 2019; Austin et al., 2021) and GSM8K-Python (Cobbe et al., 2021; Chowdhery et al., 2023; Wang et al., 2023f).
@@
Many early works argue that simply treating program synthesis as a text generation task does not utilize the underlying syntax of programming languages, and thus often use syntax-enhanced decoders to inject the target syntax as prior knowledge (Yin & Neubig, 2017). LLMs, however, have demonstrated that pretrained language models are capable of generating syntactically correct programs without loss of generality. Under this setting, researchers start to execute the generated programs and provide feedback to the generation model to inject the prior knowledge of code instead. This has recently led to the popularity of interactive coding, which we discuss in more detail in Section 7.1.
2.4 Repository-Level Coding
@@
Most tasks discussed in Section 2.1 are limited to a single file or even a single function, as cross-file code modeling poses challenges that are beyond the capability of most existing language models. Recently, however, position interpolation techniques (Chen et al., 2023b; Rozière et al., 2023; Peng et al., 2023a) have extended the context window of LLMs to hundreds of thousands of tokens, making it possible to contextualize coding activities within entire repositories. Several works have studied code completion (Shrivastava et al., 2023b; Ding et al., 2022b; Zhang et al., 2023b; Shrivastava et al., 2023a; Phan et al., 2024; Wu et al., 2024) and generation (Liao et al., 2023; Zan et al., 2024) leveraging repository-level context, and corresponding benchmarks have been proposed Liu et al. (2023j); Ding et al. (2023); Li et al. (2024a). Bairi et al. (2023) investigate the more challenging tasks of repository-level API migration and temporal editing, while Jimenez et al. (2023) introduce a related benchmark, SWE-bench.
3 Language Modeling Preliminaries
As code is ultimately a subset of natural languages, language models have been extensively used to tackle the tasks listed in Section 2. Before diving into the language models themselves, we first briefly review the preliminaries of Transformer-based language modeling in this section following the common choices of training objectives, and also some implementation designs.
3.1 Causal Language Modeling
Unidirectional language models (also known as causal language models333The training objective of such language models is Causal Language Modeling (CLM), but also referred to as Next Token Prediction.) factor the probability of a sentence into the product of each token’s conditional probability with the chain rule. A piece of input text consisting of tokens is modeled as
3.2 Masked Language Modeling
While this objective requires the model to have a deep understanding of the input text to reconstruct it, it suffers from low training efficiency, since only a small set of tokens (usually 15%) are masked (and thus “trained on”). To address this issue, Clark et al. (2020) propose ELECTRA, which is trained to discriminate whether or not each token in the input has been replaced by a BERT-like model instead, thereby computing loss on all input tokens.
3.3 Denoising Objectives
The vanilla Transformer encoder-decoder architecture combines the respective advantages of GPT and BERT. T5 (Raffel et al., 2020) is such a model pretrained with span corruption, which can be regarded as a variant of MLM. During pretraining, spans of text in the input are replaced with sentinel tokens, which play the same role as [MASK] in BERT. The noisy input is first processed by the encoder with bidirectional attention, and the masked spans are then generated autoregressively by the decoder. Formally, if spans are sampled for corruption in input , the noisy input is then constructed by replacing each span with a special token <extra_id_i>, for , and the target is constructed by concatenating all spans prepended with corresponding sentinels: [<extra_id_1>, span1, , <extra_id_k>, spank]. The model is then trained with a standard seq2seq objective, by maximizing
Lester et al. (2021) show that models pretrained with such objectives can be adapted for autoregressive language modeling with extra pretraining using the prefix language modeling objective, i.e. splitting the text into two parts, processing the first part with encoder and generating the second part with decoder.
Tay et al. (2023b) argue that span corruption is also closely related to CLM, since one can mask out the whole input text as a single span and train the decoder to generate it autoregressively. Inspired by this relation, they propose UL2, which is the combination of many span corruption objectives that differ in corruption rate and span length. Applying it to both encoder-decoder models and decoder-only models, they find that encoder-decoder models perform better under the same computation budget constraint. Other researches have also found that such encoder-decoder models generally perform better than causal decoder-only models (Wang et al., 2022b; Soltan et al., 2022).
3.4 Auxiliary Objectives
3.5 Implementation Design
As self-attention alone cannot capture the position information of each input token, the vanilla Transformer adds a non-learnable vector that’s dependent on the absolute position in the input sequence to the embedding of each token, which is known as cosine position embedding. Later works such as GPT and BERT use learnable embeddings instead, while T5 adopts relative position embedding (Shaw et al., 2018), where the information of relative position between query and key is injected in the computation of self-attention instead. A more recent scheme RoPE (Su et al., 2024) multiplies the keys and queries by a position-dependent rotation matrix, and is later shown to enable position interpolation for processing of longer sequences (Chen et al., 2023b; Rozière et al., 2023; Peng et al., 2023a). Alternatively, Press et al. (2022) propose ALiBi, which directly attenuates the attention scores according to the relative position between key and query. RoPE is popularized by its application in PaLM (Chowdhery et al., 2023) and GPT-NeoX (Black et al., 2022), while ALiBi is adopted by BLOOM (Scao et al., 2022). We refer to Zhao et al. (2023) for a survey on position embedding and interpolation in Transformers.
Apart from position embeddings, another issue in Transformer that has long troubled researchers is the fact that the complexity of self-attention scales quadratically with the input sequence length. Some works such as Reformer (Kitaev et al., 2020), Linformer (Wang et al., 2020c), Performer (Choromanski et al., 2021) and cosFormer (Qin et al., 2022b) use approximate attention to reduce this complexity, but they mostly come at the cost of degraded performance (Tay et al., 2023a; Lin et al., 2022). Other works tackle this issue from an engineering point-of-view. MQA (Shazeer, 2019) shares the same set of keys and values across all attention heads to optimize memory-to-arithmetic ratio and significantly improves inference speed at small costs of model performance. Its variant Grouped-Query Attention (GQA, Ainslie et al., 2023) takes a middle-ground approach by dividing attention heads into groups and sharing the same set of keys/values within each group. Orthogonally, Dao et al. (2022) introduce FlashAttention, which is an exact but improved implementation of self-attention that optimizes IO operations on the accelerating device via tiling to improve memory efficiency.
Another important technique for improving LLMs’ training efficiency is mixed precision training (Micikevicius et al., 2018), which stores model weights and activations in lower precision to save memory consumption. Early works use FP16 format for this low precision, while BF16 has now become more popular (Chowdhery et al., 2023; Scao et al., 2022), as it can represent the same range of floating point numbers as FP32, thus preventing numerical overflowing and underflowing during training.
@@ 4 General Language Models for Code
Since language models scaled to hundreds of billions of parameters (Brown et al., 2020; Chowdhery et al., 2023), many of them have demonstrated non-trivial coding capability, even if they are not specifically designed or trained for code. Pioneered by Codex, researchers have also found continual pretraining on code to significantly benefit language models’ performance on code555While some works refer to this process as “finetuning on code”, it is still self-supervised in nature. Thus we choose to adopt the term “extra/additional/continual pretraining” in this work to avoid confusion with supervised in-task finetuning or instruction finetuning..
@@ 4.1 Off-the-Shelf Language Models
Size HumanEval (0) MBPP (3) k=1 k=100 k=1 k=80 GPT-J1 6B 11.6 27.7 LaMDA23 137B 14.0 47.3 14.8 62.4 PaLM2 540B 26.2 76.2 36.8 75.0 GPT-NeoX4 20B 15.4 41.2 BLOOM4 176B 15.5 55.5 LLaMA5 65B 23.7 79.3 37.7 76.8 GPT-4 67.06/827 LLaMA 28 70B 29.9 89.0 45.0 81.5 Phi-1.59 1.3B 41.4 43.5 Baichuan 210 13B 17.1 30.2 Qwen11 14B 32.3 40.8 Mistral12 7B 30.5 47.5 Gemini13 Ultra 74.7 DeepSeek14 67B 42.7 57.4 Mixtral15 8x7B 40.2 60.7 DeepSeekMoE16 16B 26.8 39.2 Codex1 12B 28.8 72.3 PaLM-Coder2 540B 36.0 88.4 47.0 80.8 PaLM 2*17 S 37.6 88.4 50.0 86.8 Code LLaMA18 34B 53.7 94.7 56.2 Code-Qwen11 14B 45.1 51.4
Closed-source models, on the other hand, perform generally better. LaMDA (Thoppilan et al., 2022) and PaLM (Chowdhery et al., 2023), whose pretraining dataset contains 12.5% and 5% code respectively, achieve 14.0 and 26.2 pass@ performance on HumanEval, while GPT-4 (OpenAI, 2023) set a staggering record of 67.0 (and an early version is reported by Bubeck et al. (2023) to be 82) that until recently has remained higher than any specialized models pretrained or instruction-finetuned for code.
More recently, the general trend has been to train smaller models with larger datasets, following the revised scaling law (Hoffmann et al., 2022). Baichuan 2 (Yang et al., 2023a), for example, is a 13B model trained on 2.6T tokens, while Qwen (Bai et al., 2023) is a 14B model trained on 3T tokens. They achieve 17.1 and 32.3 pass@1 on HumanEval, respectively. Li et al. (2023i), however, demonstrate that models as small as 1.3B can acquire coding capability that’s comparable to much larger models while also maintaining a reasonable performance on general text processing and even manifesting some emergent abilities (Wei et al., 2022c) such as chain-of-though reasoning (Wei et al., 2022d). Their model, Phi-1.5, is trained on 21B tokens of textbook data generated by ChatGPT, and 100B tokens of filtered web data from Stack Overflow and Refined Web (Penedo et al., 2023), and attains 41.4 pass@ performance on HumanEval.
Another rising trend is mixture-of-expert models (Lepikhin et al., 2021; Fedus et al., 2022; Du et al., 2022). Notably, Jiang et al. (2024a) recently introduce Mixtral 8x7B, where only 13B parameters are activated for each token during inference, but achieving 40.2 on HumanEval and surpassing much larger dense models such as LLaMA 2. Similarly, Dai et al. (2024) present DeepSeekMoE, where only 2.8B parameters are activated in a 16B model, scoring 26.8 on HumanEval.
The exact performance of these models are presented in Table 1.